4. What is the annual interest rate if a simple interest loan of $15,000 for four months yields $975 interest? (20 pts.)
Answer #1

Ans: 19.5

Explanation:

Total Interest earned in 4 months = $975

Total Interest earned in 1 months = $975 / 4 = $243.75

Total Interest earned in 12 months ( annual ) = $243.75 * 12 = $2925

So annual interest rate = ( 2925 / 15000 ) * 100

= 0.195 * 100

= 19.5